# TTPXHunter: Actionable Threat Intelligence Extraction as TTPs form Finished Cyber Threat Reports

### 05 Mar 2024 v1

@@ VII Conclusion
This research demonstrates the efficiency of domain-specific language models in extracting threat intelligence in terms of TTPs, which can share the attack patterns and accelerate the threat response and detection mechanisms. The tool TTPXHunter extends the TTPHunter’s capability over multiple dimensions, such as expanding to the full spectrum of TTP extraction and improving efficiency by leveraging domain-specific language models. We evaluate the efficiency of TTPXHunter over the prepared augmented sentence-TTP dataset and report-TTP dataset. On the augmented dataset, TTPXHunter outperforms both BERT-variant models, i.e., TRAM and TTPHunter. TTPXHunter also outperforms the state-of-the-art TTP extraction methods by achieving the highest F1 score and lowest hamming loss. TTPXHunter’s performance over the report dataset demonstrates the model’s efficiency in capturing relevant sentences from threat reports and correctly classifying them to the TTP class. The conversion of extracted TTPs to STIX makes integrating threat intelligence into security operations easier. TTPXHunter aids in improving the threat analyst’s capability to share intelligence, analyze threats, understand the modus operandi of sophisticated threat actors, and emulate their behavior for red teaming. Therefore, TTPXHunter can support various cybersecurity teams, including red, blue, and purple teams in the organization.

### 21 Mar 2024 v3

@@ VII Conclusion
This research demonstrates the efficiency of domain-specific language models in extracting threat intelligence in terms of TTPs, which can share the attack patterns and accelerate the threat response and detection mechanisms. The tool TTPXHunter extends the TTPHunter’s capability over multiple dimensions, such as expanding to the full spectrum of TTP extraction and improving efficiency by leveraging domain-specific language models. We evaluate the efficiency of TTPXHunter over the prepared augmented sentence-TTP dataset and report-TTP dataset. On the augmented dataset, TTPXHunter outperforms both BERT-variant models, i.e., TRAM and TTPHunter. TTPXHunter also outperforms the state-of-the-art TTP extraction methods by achieving the highest F1 score and lowest hamming loss. TTPXHunter’s performance over the report dataset demonstrates the model’s efficiency in capturing relevant sentences from threat reports and correctly classifying them to the TTP class. The conversion of extracted TTPs to STIX makes integrating threat intelligence into security operations easier. TTPXHunter aids in improving the threat analyst’s capability to share intelligence, analyze threats, understand the modus operandi of sophisticated threat actors, and emulate their behavior for red teaming. Therefore, TTPXHunter can support various cybersecurity teams, including red, blue, and purple teams in an organization.